A neutron star of mass 4.0 x 1030kg has a radius of about 12km. Calculate its mean density . comment on the answer


1
Expert's answer
2021-02-15T00:54:59-0500

Let's find the mean density of the neutron star:


"\rho=\dfrac{m}{V}=\dfrac{m}{\dfrac{4}{3}\pi r^3},""\rho=\dfrac{4.0\cdot10^{30}\ kg}{\dfrac{4}{3}\pi\cdot(1.2\cdot10^4\ m)^3}=5.53\cdot10^{17}\ \dfrac{kg}{m^3}."

As we can see from calculations, the neutron star has an incredibly high density comparable to or greater than the density of an atomic nucleus.